# Survey on the Evaluation of Generative Models in Music

### v1

@@ 2.1.4. Conditioning
Thus, this direction of evaluation —focusing on individual output properties and whether they match the user input or training dataset properties— is mostly useful for the verification that the training process was successful. The evaluation of individual properties allows a very targeted quantitative validation of the semantic relevance of the output with respect to certain, usually narrowly defined characteristics. However, measuring and interpreting specific characteristics of the output as a proxy for an overall assessment of the output is questionable at best as any set of characteristics represents only a small subset of a possibly infinite number of musical characteristics, and thus can only give a snapshot of one facet of the generated output. It is particularly problematic if the measured property was explicitly used as a training target or loss function, cf. Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure” (Strathern, 1997; Goodhart, 2015).
@@ 2.2. Usability & user experience
Outputs from generative music systems are used and appreciated by people from musicians to audiences. In this section, we focus on evaluation of people’s direct interaction with generative systems. Viewing this through a HCI (Preece et al., 2015) lens we refer to the people as users who interact with the generative system through a real-time user interface. Unlike evaluation of system outputs using listening tests described above, evaluation here is concerned with understanding the interaction between the user and the generative model — the human-in-the-loop. This aligns with recent Human-Centered AI discourse (Shneiderman, 2022; Ozmen Garibay et al., 2023) advocating for the use of HCI methods to evaluate and inform the design of AI systems which balance automation with human control. There are two main aspects of interacting with computers, and AI models specifically, that are usually evaluated, the usability and the user experience.
@@ 3.1.1. Turing tests
Turing-type tests —where a listener must identify whether a musical selection was made by a human or AI— remain one of the most common forms of subjective evaluation (Hernandez-Olivan et al., 2022). In its most basic form, the Turing test is a useful metric in the sense that the method is simple (i.e., typically binary forced selection), and that it offers a theoretically unbiased subjective evaluation by, in principle, implicitly evaluating a model’s output according to a scale of “humanness.” If a listener cannot tell apart machine from human-generated musical output, then the implication is that the machine is “at least as good” as a human. However, this logic only follows under the right conditions, which may not be met in small-scale, ad-hoc experiments. For example, what makes a Bach chorale different from another piece of music from the classical genre may be non-evident to a lay listener; similar arguments can be made for a jazz solo. In other words, recognition of the norms of a particular musical style typically takes at least a small degree of expertise (Pearce, 2015). Another consideration is the material used in the test itself. Since not all of the output material can be evaluated, only a small subset of the model output is used in the test. However, depending on how this material is selected, this selection may not be representative of the model output overall. In addition, inferential statistical tests designed to test an alternative hypothesis against a null hypothesis, such as t-tests, are commonly used to evaluate the outcome of a Turing test. Yet typically the desired outcome is that the null hypothesis (i.e., no difference) is actually supported. This inappropriate use of such a test will “bias” towards supporting a null hypothesis is compounded by studies that rely on small sample sizes or exhibit high variance. The use of Turing tests in evaluating AI output has been criticized for a variety of reasons, most notably for its lack of sophistication, and for the tool being repurposed for something other than what it was intended for (Hernández-Orallo, 2020), which was as an evaluation of intelligence and not of aesthetics. It is important to note is that the Turing test conflates a lack of distinction (between human and computer) with aesthetic or creative success. It is easily conceivable that a model trained on student input, for instance, would ask a listener to disambiguate between two samples both representing ‘amateur.’ In this case, a lack of ability to discriminate between the samples does not conclude that the output is of high aesthetic or creative value. Finally, an inherent problem with using a Turing test as a subjective evaluation metric is that it inherently over-rewards imitation over creativity (Pease and Colton, 2011). Nevertheless, if used appropriately, the Turing test can be used to assess model output, provided that these considerations are taken into account and that the test is primarily used to support a conformity to a baseline standard of a specific musical style, rather than an indication of musically or aesthetically valued output.
@@ 3.2.2. Aesthetic and musical qualities
Metrics targeting musical qualities are commonly distribution-based, meaning that the similarity between a generated and a target distribution is measured for a specific musical descriptor or property. Typically, these descriptors are low-level representations of music (Ji et al., 2023) and some of these features and distributions might be more musically relevant than others. For example, pitch distributions are commonly applied to MIDI numbers or pitch values (e.g., A4) as opposed to much more meaningful pitch or chord distributions that take key information into account, such as scale degree (the position of a pitch class in a tonal context) or Roman numeral (the function of a chord given its context). Still, basic features such as pitch difference and the ratio of in-key pitches continue to be employed in recent works and are sometimes referred to as “music theory evaluation” (Guo et al., 2023). Over time, many descriptors for musical qualities have been introduced (Meredith, 2016); Yang and Lerch (2020) proposed a set of simple metrics separately targeting tonal and rhythmic content. Similarly, Garcia-Valencia et al. (2021) utilized the concepts proposed by Tymoczko (2011) to assess qualities like the smoothness of melody transitions, pitch diversity, and the occurrence of local notes to provide an analysis of melodic structure with rhythmic features. Dervakos et al. (2021) introduced a framework based on the basic consonance aspects of melodies that allows the construction of a variety of metrics. They used this framework to construct four heuristic properties, polyphonicity, used-pitch-classes per bar, total number of pitches, and the total number of pitch classes, to address fundamental musical properties that are reliable and interpretable. These examples show the variety of descriptors that have been used to model musical properties. Table 1 provides an overview of common descriptors proposed in the literature.
@@
Furthermore, scholars have pointed out the importance of higher level features such as form and repetition in contributing to the improvement of generative music systems (e.g., (Dai et al., 2022)). However, specific metrics to measure these higher-level features have not yet been proposed.
@@ 3.2.3. Audio quality
For any generative audio system, assessing the audio quality of the outputs is integral to understand whether the generated signals would can be considered by humans as high-quality signals. Simple error measurements such as the MSE, the Mean Absolute Error (MAE), or the SNR have been long shown to be ineffective as “quality” measurements (Karjalainen, 1985), so that listening studies are often considered to be the ultimate way to assess audio quality. Regardless, objective metrics for measuring audio quality have been proposed.
@@
Many of the contemporary approaches to the problem use the aforementioned distribution comparison metrics. By and large, these methods extract neural representations or embeddings for both the target and testing sets. The primary argument for the usage of distribution comparison methods hinges on the notion that minimal divergence between the two distributions is indicative of the two sets being roughly identical in terms of quality. The most popular distribution divergence metric used in evaluating generative quality is the Fréchet Audio Distance (FAD) (Kilgour et al., 2019), which computes the difference between the two embedding distributions using a Wasserstein distance on VGGish embeddings (Hershey et al., 2017). The perceptual relevance of FAD has, however, been questioned (Vinay and Lerch, 2022). Gui et al. (2024) propose a variation of the FAD for audio-based music quality measurement called FAD. FAD is an extension of FAD that mitigates the sample-size bias by estimating the behavior of the metric as if it had an infinite number of samples by using Quasi Monte-Carlo integrals. Gui et al. (2024) investigate replacing the VGGish embedding with different embeddings and report that —given the right embeddings and mitigating sample size bias— FAD can be a useful indicator of quality, a result that is intended to assuage concerns raised by previous results (Vinay and Lerch, 2022). An alternative to FAD is to use Maximum-Mean Discrepancy (MMD) (Binkowski et al., 2018; Nistal et al., 2020), which is a two-sample test computed over the kernel embedding of the training and test data. For more details on the mechanics of MMD, we refer the reader to Gretton et al. (2012). Recent work in the image domain has shown that MMD with CLIP embeddings is a better approximator of generative image quality than Fréchet distances (Jayasumana et al., 2023).
@@
Fréchet Audio Distance VGGish 0 to $\infty$ $\downarrow$
@@
Other notable approaches to estimating audio quality include DNSMOS (Reddy et al., 2021) NDB/$k$ (Richardson and Weiss, 2018) and ViSQOL (Hines et al., 2015), which are predominantly speech quality focused approaches. ViSQOL and DNSMOS are speech quality estimation methods trained using results from large-scale MOS listening studies. ViSQOL is a paired metric that uses spectrogram patches and DNSMOS is a no-reference metric that uses a model trained to predict what a human might rate a signal. NDB/$k$ is a metric originally proposed for images that has been used to evaluate some generative systems such as Diffwave (Kong et al., 2020). It clusters the features of the training set using k-means clustering and Voronoi cell partitioning. To compute the metric, the number of statistically different bins or cells are computed. Recent work has shown that in speech, similarity is not a reliable proxy for quality and no-reference metrics, such as the ones mentioned above, might be better at estimating quality (Manocha et al., 2022).
@@ 3.2.4. Originality
Sturm et al. (2019b) note that for music —unlike for text— automated plagiarism detection does not reliably work, and worse, no clear standards exist about what type and amount of alterations make a piece of music novel as opposed to plagiarized. Thus, approaches such as the Authenticity metric proposed by Alaa et al. (2022), estimating the probability of a generated sample being copied from the training data, are only of limited use. Still, approaches to objectively measure plagiarism have been proposed, usually framed as a music similarity task. Most of the seminal work can be grouped into (i) melodic similarity measures based on symbolic input and (ii) general audio-based similarity measures. To measure symbolic melodic similarity, a variety of metrics have been proposed. Most of them are based on some form of sequence similarity (Park et al., 2005; Müllensiefen and Pendzich, 2009; Wolf and Müllensiefen, 2011; Cason and Müllensiefen, 2012; De Prisco et al., 2017a, b) or vector-based similarity measures hand-designed (Malandrino et al., 2022) or trained as end-to-end systems (Park et al., 2022). Some of these algorithms have been tested against real-world court decisions, however, the test set sizes are necessarily small and external validity is hard to verify. Audio-based similarity (Lerch, 2023) is a multi-dimensional problem at the risk of confounding dimensions of score similarity (melody, harmony, rhythm, etc.) with performance similarity (tempo, playing techniques, etc., but particularly timbre). Systems for audio-based plagiarism detection have been proposed to use Non-Negative Matrix Factorization to decompose the audio (De et al., 2012), MFCC vector-based representations of audio (Suneja and Bansal, 2015), or similarity measures inspired by audio fingerprinting approaches (Borkar et al., 2021; López-García et al., 2022). Commonly, however, plagiarism is understood in terms of score similarity (with the very prominent exception of sampling, where audio is copied and mixed into a new musical artifact (Gururani and Lerch, 2017)). A study into the contributing factors of court cases on plagiarism has been presented by Yuan et al. (2023).
@@
The evaluation of creativity poses an unsolved problem. Jordanous (2012) notes the longstanding lack of attention to evaluation and lack of evaluation standards in the computational creativity community due to “difficulties in defining what it means for a computer to be creative; indeed, there is no consensus on this for human creativity, let alone its computational equivalent.” Indeed, in a survey of 75 papers on computational creativity, the author found that in one third of the papers, ‘creativity’ was not mentioned, and that only one third of the papers actually attempted to evaluate the creativity of their systems. Moreover, “[o]ccurrences of creativity evaluation by people outside the system implementation team were rare.” Judging from the scarcity of proposed evaluation standards since the publication of this paper —despite the fact that it has been over a decade since its publication and that creative AI is presently in its heyday— it seems that this issue continues to persist. As mentioned above, one can distinguish between the evaluation of a computational system itself, and such a system’s output (i.e., “the product/process debate” (Jordanous, 2012)). However, since the output is restricted or constrained by the system, it can be helpful to conceptualize a combined methodology for the evaluation of both the system (in terms of its creative potential) and its output. Jordanous (2012) presents such a framework, referred to as “SPECS” (Standardised Procedure for Evaluating Creative Systems). The SPECS method —which is not only well recommended (O’Neill and Loughran, 2017), but unlike other proposed models has actually been used in practice (Ostermann et al., 2021; Plucker and Makel, 2010)— asks the creator/evaluator to adhere to a specific working definition of creativity that includes two or more “components” that are evaluated independently according to some standard, such as skill, novelty, value, etc., and compared against an appropriate system/output. These evaluations may be quantitative or qualitative, however, in this case, both still involve significant investment of time and resources into evaluation with human subjects. While other models and frameworks have been proposed for attempting to theoretically quantitatively assess creative output, such at the FACE and IDEA models (Pease and Colton, 2011), to our knowledge, there remain no adopted standardized practices or procedures in place for assessing creativity without human intervention at this time.

### v2

@@ 2.1.4. Conditioning
Thus, this direction of evaluation —focusing on individual output properties and whether they match the user input or training dataset properties— is mostly useful for the verification that the training process was successful. The evaluation of individual properties allows a very targeted quantitative validation of the semantic relevance of the output with respect to certain, usually narrowly defined characteristics. However, measuring and interpreting specific characteristics of the output as proxies for an overall assessment of the output is, at best, questionable. Any set of characteristics represents only a small subset of a possibly infinite number of musical characteristics, and thus can only give a snapshot of one facet of the generated output. It is particularly problematic if the measured property was explicitly used as a training target or loss function, cf. Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure” (Strathern, 1997; Goodhart, 2015).
A special, but recently very popular form of conditioning is the text prompt. Unlike the evaluation of other conditionings, measuring the (perceptual) alignment of the text prompt with the musical output is challenging due to the undefined structure and terminology or the prompt and the multitude of potentially impacted musical properties. Given these challenges, most evaluation strategies focus on global measures of fit (Huang et al., 2025; Grötschla et al., 2025; Liu et al., 2025).
@@ 2.2. Usability & user experience
Outputs from generative music systems are used and appreciated by people from musicians to audiences. In this section, we focus on evaluation of people’s direct interaction with generative systems. Viewing this through a HCI (Preece et al., 2015) lens we refer to the people as users who interact with the generative system through a real-time user interface. Unlike evaluation of system outputs using listening tests described above, evaluation here is concerned with understanding the interaction between the user and the generative model — the human-in-the-loop. This aligns with recent Human-Centered AI discourse (Shneiderman, 2022; Ozmen Garibay et al., 2023) advocating for the use of HCI methods to evaluate and inform the design of AI systems which balance automation with human control. There are two main aspects of interacting with computers, and AI models specifically, that are usually evaluated: the usability and the user experience.
@@ 3.1.1. Turing tests
Turing-type tests —where a listener must identify whether a musical selection was made by a human or AI— remain one of the most common forms of subjective evaluation (Hernandez-Olivan et al., 2022). In its most basic form, the Turing test is a useful metric in the sense that the method is simple (i.e., typically binary forced selection), and that it offers a theoretically unbiased subjective evaluation by, in principle, implicitly evaluating a model’s output according to a scale of “humanness.” If a listener cannot tell apart machine from human-generated musical output, then the implication is that the machine is “at least as good” as a human. However, this logic only follows under the right conditions, which may not be met in small-scale, ad-hoc experiments. For example, what makes a Bach chorale different from another piece of music from the classical genre may be non-evident to a lay listener; similar arguments can be made for a jazz solo. In other words, recognition of the norms of a particular musical style typically takes at least a small degree of expertise (Pearce, 2015). Another consideration is the material used in the test itself. Since not all of the output material can be evaluated, only a small subset of the model output is used in the test. However, depending on how this material is selected, this selection may not adequately represent the model output overall. In addition, inferential statistical tests designed to test an alternative hypothesis against a null hypothesis, such as t-tests, are commonly used to evaluate the outcome of a Turing test. Yet typically the desired outcome is that the null hypothesis (i.e., no difference) is actually supported. This inappropriate use of such a test will “bias” towards supporting a null hypothesis is compounded by studies that rely on small sample sizes or exhibit high variance. The use of Turing tests in evaluating AI output has been criticized for a variety of reasons, most notably for its lack of sophistication, and for the tool being repurposed for something other than what it was intended for (Hernández-Orallo, 2020), which was as an evaluation of intelligence and not of aesthetics. It is important to note is that the Turing test conflates indistinguishability (between human and computer) with aesthetic or creative success. It is easily conceivable that a model trained on student input, for instance, would ask a listener to disambiguate between two samples both representing ‘amateur.’ In this case, a lack of ability to discriminate between the samples does not conclude that the output is of high aesthetic or creative value. Finally, an inherent problem with using a Turing test as a subjective evaluation metric is that it inherently over-rewards imitation over creativity (Pease and Colton, 2011). Nevertheless, if used appropriately, the Turing test can be used to assess model output, provided that these considerations are taken into account and that the test is primarily used to support a conformity to a baseline standard of a specific musical style, rather than an indication of musically or aesthetically valued output.
@@ 3.2.2. Aesthetic and musical qualities
Metrics targeting musical qualities are commonly distribution-based, meaning that the similarity between a generated and a target distribution is measured for a specific musical descriptor or property. Typically, these descriptors are low-level representations of music (Ji et al., 2023) and some of these features and distributions might be more musically relevant than others. For example, pitch distributions are commonly applied to MIDI numbers or pitch values (e.g., A4) as opposed to much more meaningful pitch or chord distributions that take key information into account, such as scale degree (the position of a pitch class in a tonal context) or Roman numeral (the function of a chord given its context). Still, basic features such as pitch difference and the ratio of in-key pitches or entropy of pitch and chords (Wang et al., 2023) continue to be employed in recent works and are sometimes referred to as “music theory evaluation” (Guo et al., 2023). Over time, many descriptors for musical qualities have been introduced (Meredith, 2016); Yang and Lerch (2020) proposed a set of simple metrics separately targeting tonal and rhythmic content. Similarly, Garcia-Valencia et al. (2021) utilized the concepts proposed by Tymoczko (2011) to assess qualities like the smoothness of melody transitions, pitch diversity, and the occurrence of local notes to provide an analysis of melodic structure with rhythmic features. Dervakos et al. (2021) introduced a framework based on the basic consonance aspects of melodies that allows the construction of a variety of metrics. They used this framework to construct four heuristic properties, polyphonicity, used-pitch-classes per bar, total number of pitches, and the total number of pitch classes, to address fundamental musical properties that are reliable and interpretable. These examples show the variety of descriptors that have been used to model musical properties. Table 1 provides an overview of common descriptors proposed in the literature.
@@
Furthermore, scholars have pointed out the importance of higher level features such as form and repetition in contributing to the improvement of generative music systems (e.g., (Dai et al., 2022)). However, specific metrics to measure these higher-level features have not yet been proposed. Instead, metrics such as the Fréchet Distance (see Sect. 3.2.3) utilizing trained embeddings with opaque or unknown musical meaning (VGGish (Hershey et al., 2017), CLAP (Wu et al., 2023), etc.) have been increasingly applied to music evaluation (Gui et al., 2024; Retkowski et al., 2025; Huang et al., 2025; Chung et al., 2025).
@@ 3.2.3. Audio quality
For any generative audio system, assessing the audio quality of the outputs is integral to understand whether the generated signals can be considered high-quality by human listeners. Simple error measurements such as the MSE, the Mean Absolute Error (MAE), or the SNR have been long shown to be ineffective as “quality” measurements (Karjalainen, 1985), so that listening studies are often considered to be the ultimate way to assess audio quality. Regardless, objective metrics for measuring audio quality have been proposed.
@@
Many of the contemporary approaches to the problem use the aforementioned distribution comparison metrics. By and large, these methods extract neural representations or embeddings for both the target and testing sets. The primary argument for the usage of distribution comparison methods hinges on the notion that minimal divergence between the two distributions is indicative of the two sets being roughly identical in terms of quality. The most popular distribution divergence metric used in evaluating generative quality is the Fréchet Audio Distance (FAD) (Kilgour et al., 2019), which computes the difference between the two embedding distributions using a Wasserstein distance on VGGish embeddings (Hershey et al., 2017), and more recently on a variety of other embeddings, such as PANNs (Kong et al., 2020a), CLAP (Wu et al., 2023), and Encodec (Défossez et al., 2023) (cf. (Gui et al., 2024; Grötschla et al., 2025)). The perceptual relevance of FAD has, however, been questioned for both audio quality evaluation (Vinay and Lerch, 2022) and for music evaluation (Huang et al., 2025). Gui et al. (2024) propose a variation of the FAD for audio-based music quality measurement called FAD. FAD is an extension of FAD that mitigates the sample-size bias by estimating the behavior of the metric as if it had an infinite number of samples by using Quasi Monte-Carlo integrals. Gui et al. (2024) report that —given the right embeddings and mitigating sample size bias— FAD can be a useful indicator of quality, a result that is intended to assuage concerns raised by previous results (Vinay and Lerch, 2022). An alternative to FAD is to use Maximum-Mean Discrepancy (MMD) (Binkowski et al., 2018; Nistal et al., 2020), which is a two-sample test computed over the kernel embedding of the training and test data. For more details on the mechanics of MMD, we refer the reader to Gretton et al. (2012). Recent work in the image domain has shown that MMD with CLIP embeddings is a better approximator of generative image quality than Fréchet distances (Jayasumana et al., 2023). Chung et al. (2025) propose to adapt the MMD to the Kernel Audio Distance (KAD) to replace FAD.
@@
Fréchet Audio Distance VGGish, CLAP, etc. 0 to $\infty$ $\downarrow$
@@
Other notable approaches to estimating audio quality include DNSMOS (Reddy et al., 2021) NDB/$k$ (Richardson and Weiss, 2018) and ViSQOL (Hines et al., 2015), which are predominantly speech quality focused approaches. ViSQOL and DNSMOS are speech quality estimation methods trained using results from large-scale MOS listening studies. ViSQOL is a paired metric that uses spectrogram patches and DNSMOS is a no-reference metric that uses a model trained to predict what a human might rate a signal. NDB/$k$ is a metric originally proposed for images that has been used to evaluate some generative systems such as Diffwave (Kong et al., 2020b). It clusters the features of the training set using k-means clustering and Voronoi cell partitioning. To compute the metric, the number of statistically different bins or cells are computed. Recent work has shown that in speech, similarity is not a reliable proxy for quality and no-reference metrics, such as the ones mentioned above, might be better at estimating quality (Manocha et al., 2022).
@@ 3.2.4. Originality
Sturm et al. (2019b) note that for music —unlike for text— automated plagiarism detection does not reliably work, and worse, no clear standards exist about what type and amount of alterations make a piece of music novel as opposed to plagiarized. Thus, approaches such as the Authenticity metric proposed by Alaa et al. (2022), estimating the probability of a generated sample being copied from the training data, are only of limited use. Still, approaches to objectively measure plagiarism have been proposed, usually framed as a music similarity task. Most of the seminal work can be grouped into (i) melodic similarity measures based on symbolic input and (ii) general audio-based similarity measures. To measure symbolic melodic similarity, a variety of metrics have been proposed. Most of them are based on some form of sequence similarity (Park et al., 2005; Müllensiefen and Pendzich, 2009; Wolf and Müllensiefen, 2011; Cason and Müllensiefen, 2012; De Prisco et al., 2017a, b) or vector-based similarity measures hand-designed (Malandrino et al., 2022) or trained as end-to-end systems (Park et al., 2022; Lu et al., 2025). Some of these algorithms have been tested against real-world court decisions, however, the test set sizes are necessarily small and external validity is hard to verify. Audio-based similarity (Lerch, 2023) is a multi-dimensional problem at the risk of confounding dimensions of score similarity (melody, harmony, rhythm, etc.) with performance similarity (tempo, playing techniques, etc., but particularly timbre). Systems for audio-based plagiarism detection have been proposed to use Non-Negative Matrix Factorization to decompose the audio (De et al., 2012), MFCC vector-based representations of audio (Suneja and Bansal, 2015), similarity measures inspired by audio fingerprinting approaches (Borkar et al., 2021; López-García et al., 2022), or utilizing pre-trained embeddings for similarity measurement (Batlle-Roca et al., 2024). Commonly, however, plagiarism is understood in terms of score similarity (with the very prominent exception of sampling, where audio is copied and mixed into a new musical artifact (Gururani and Lerch, 2017)). A study into the contributing factors of court cases on plagiarism has been presented by Yuan et al. (2023).
@@
The evaluation of creativity poses an unsolved problem. Jordanous (2012) notes the longstanding lack of attention to evaluation and lack of evaluation standards in the computational creativity community due to “difficulties in defining what it means for a computer to be creative; indeed, there is no consensus on this for human creativity, let alone its computational equivalent.” Indeed, in a survey of 75 papers on computational creativity, the author found that in one third of the papers, ‘creativity’ was not mentioned, and that only one third of the papers actually attempted to evaluate the creativity of their systems. Moreover, “[o]ccurrences of creativity evaluation by people outside the system implementation team were rare.” Judging from the scarcity of proposed evaluation standards since the publication of this paper —despite the fact that it has been over a decade since its publication and that creative AI is presently in its heyday— it seems that this issue continues to persist. As mentioned above, one can distinguish between the evaluation of a computational system itself, and such a system’s output (i.e., “the product/process debate” (Jordanous, 2012)). However, since the output is restricted or constrained by the system, it can be helpful to conceptualize a combined methodology for the evaluation of both the system (in terms of its creative potential) and its output. Jordanous (2012) presents such a framework, referred to as “SPECS” (Standardised Procedure for Evaluating Creative Systems). The SPECS method —which is not only well recommended (O’Neill and Loughran, 2017), but unlike other proposed models has actually been used in practice (Ostermann et al., 2021; Plucker and Makel, 2010)— asks the creator/evaluator to adhere to a specific working definition of creativity that includes two or more “components” that are evaluated independently according to some standard, such as skill, novelty, value, etc., and compared against an appropriate system/output. These evaluations may be quantitative or qualitative, however, in this case, both still involve significant investment of time and resources into evaluation with human subjects. While other models and frameworks have been proposed for attempting to theoretically quantitatively assess creative output, such at the FACE and IDEA models (Pease and Colton, 2011), to our knowledge, no adopted standardized practices or procedures currently exist for assessing creativity without human intervention.